# Bounding the SNPR distance between two tree-child networks using generalised agreement forests

### v1

@@ 4. Properties of extensions
    Suppose that both of $e$ and $e^{\prime}$ are contained in one of ${\mathcal{M}}$ and ${\mathcal{M}}^{\prime}$, and neither $e$ nor $e^{\prime}$ is contained in the other embedding. Without loss of generality, we may assume that both of $e$ and $e^{\prime}$ are contained in ${\mathcal{M}}$ and, therefore, in ${\mathcal{R}}$. As ${\mathcal{M}}^{\prime}$ does not contain a vertex with out-degree zero that is not in $X\cup\{\rho\}$, it follows that $v\notin{\mathcal{M}}^{\prime}$. Thus, by Lemma 4.2(ii), exactly one of $e$ and $e^{\prime}$ is in ${\mathcal{R}}^{\prime}$.
@@
    Suppose that exactly one of $e$ and $e^{\prime}$ is contained in one of ${\mathcal{M}}$ and ${\mathcal{M}}^{\prime}$ and neither $e$ nor $e^{\prime}$ is contained in the other embedding. Without loss of generality, we may assume that ${\mathcal{M}}$ contains $e$ and does not contain $e^{\prime}$, and that ${\mathcal{M}}^{\prime}$ contains neither $e$ nor $e^{\prime}$. Evidently $e\in{\mathcal{R}}$ and, by the definition of an extension, $e^{\prime}\notin{\mathcal{R}}$. Moreover, as $v\notin{\mathcal{M}}^{\prime}$ it follows from Lemma 4.2(ii) that exactly one of $e$ and $e^{\prime}$ is in ${\mathcal{R}}^{\prime}$.
@@
    Suppose that exactly one of $e$ and $e^{\prime}$ is contained in each of ${\mathcal{M}}$ and ${\mathcal{M}}^{\prime}$. Again by Lemma 4.2(i), exactly one of $e$ and $e^{\prime}$ is contained in each of ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$.
@@
    Suppose that both of $e$ and $e^{\prime}$ are contained in one of ${\mathcal{M}}$ and ${\mathcal{M}}^{\prime}$, and exactly one of $e$ and $e^{\prime}$ is contained in the other embedding. Without loss of generality, we may assume that both of $e$ and $e^{\prime}$ are contained in ${\mathcal{M}}$ and, therefore, in ${\mathcal{R}}$. Then, by Lemma 4.2(i), exactly one of $e$ and $e^{\prime}$ is contained in ${\mathcal{R}}^{\prime}$.
@@
The last proposition motivates the following terminology. Let ${\mathcal{D}}$ be a phylogenetic digraph of a phylogenetic network ${\mathcal{N}}$ on $X$, and let ${\mathcal{R}}$ be an extension of ${\mathcal{D}}$ in ${\mathcal{N}}$. We set $c_{\mathcal{D}}=|E_{\mathcal{N}}-E_{\mathcal{R}}|$ and refer to $c_{\mathcal{D}}$ as the cut size of ${\mathcal{D}}$ in ${\mathcal{N}}$. By Proposition 4.3, $c_{\mathcal{D}}$ is well defined.
@@
Let ${\mathcal{M}}$ be the embedding of ${\mathcal{D}}$ in ${\mathcal{N}}$ that underlies ${\mathcal{R}}$. If ${\mathcal{R}}$ is not a root extension of ${\mathcal{D}}$ in ${\mathcal{N}}$, then there exists a reticulation edge $(u,v)$ in ${\mathcal{N}}$ such that $(u,v)\in{\mathcal{R}}$ and $(u,v)\notin{\mathcal{M}}$. Since ${\mathcal{N}}$ is tree-child, $u$ is a tree vertex. Let $v^{\prime}$ be the child of $u$ in ${\mathcal{N}}$ with $v\neq v^{\prime}$. Clearly, $v^{\prime}$ is a tree vertex or a leaf in ${\mathcal{N}}$, and $(u,v^{\prime})\notin{\mathcal{R}}$. Let $R_{v}$ be the component in ${\mathcal{R}}$ that contains $v$, and let $R_{v^{\prime}}$ be the component in ${\mathcal{R}}$ that contains $v^{\prime}$. If $R_{v}$ and $R_{v^{\prime}}$ are distinct, obtain $R^{\prime}_{v}$ from $R_{v}$ by deleting each vertex $s$ for which there exists a directed path from $s$ to $u$, and obtain $R^{\prime}_{v^{\prime}}$ from $R_{v^{\prime}}$ by adding $(u,v^{\prime})$ and each edge of $R_{v}$ that is contained in a directed path ending at $u$ and, if $R_{v}$ and $R_{v^{\prime}}$ are not distinct, obtain $R^{\prime}_{v}$ from $R_{v}$ by deleting $(u,v)$ and adding $(u,v^{\prime})$. Then =(${\mathcal{R}}^{\prime}=({\mathcal{R}}-\{R_{v},R_{v^{\prime}}\})\cup\{R_{v}^{\prime},R_{v^{\prime}}^{\prime}\}$ is an extension of ${\mathcal{D}}$ in ${\mathcal{N}}$ with cut size $c_{\mathcal{D}}$. It is straightforward to check that ${\mathcal{M}}$ is the embedding of ${\mathcal{D}}$ in ${\mathcal{N}}$ that underlies ${\mathcal{R}}^{\prime}$ and that ${\mathcal{R}}^{\prime}$ has one reticulation edge less than ${\mathcal{R}}$. If ${\mathcal{R}}^{\prime}$ is not a root extension of ${\mathcal{D}}$ in ${\mathcal{N}}$, then repeat the construction for a reticulation edge in ${\mathcal{N}}$ that is in ${\mathcal{R}}^{\prime}$ but not in ${\mathcal{M}}$ until no such edge remains. This completes the proof of the lemma. ∎
@@
Figure 2. A phylogenetic network ${\mathcal{N}}$, a phylogenetic digraph ${\mathcal{D}}$ of ${\mathcal{N}}$, and two root extensions ${\mathcal{R}}$ and ${\mathcal{R}}^{\prime}$ (indicated by thick lines) of ${\mathcal{D}}$ in ${\mathcal{N}}$ with $|E_{\mathcal{N}}-E_{\mathcal{R}}|=6\neq 5=|E_{\mathcal{N}}-E_{{\mathcal{R}}^{\prime}}|$.
We end this section by noting that Proposition 4.3 does not hold for two root extensions of arbitrary phylogenetic networks (see Figure 2 for an example). Nevertheless, we have the following result for tree-child networks. Its proof is similar to that of Proposition 4.3 and is omitted.
Proposition 4.5.
@@
5. Measures of dissimilarity between two tree-child networks
@@
Let ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ be two tree-child networks on $X$. Let ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ be a collection of leaf-labelled acyclic digraphs. If ${\mathcal{D}}$ is a phylogenetic digraph of ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$, we refer to ${\mathcal{D}}$ as an agreement digraph of ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$. Now let ${\mathcal{D}}$ be an agreement digraph for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$. If ${\mathcal{D}}$ is tree-child, we say that ${\mathcal{D}}$ is an agreement tree-child digraph for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$. In the remainder of this paper, we are particularly interested in agreement digraphs of ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ whose cut size is minimum. To this end, let ${\mathcal{D}}$ be an agreement tree-child digraph for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$, and let $c_{\mathcal{D}}$ and $c^{\prime}_{\mathcal{D}}$ be the cut size of ${\mathcal{D}}$ in ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$, respectively. Then ${\mathcal{D}}$ is called a maximum agreement tree-child digraph for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ if the sum $c_{\mathcal{D}}+c^{\prime}_{\mathcal{D}}$ is minimum over all agreement tree-child digraphs for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$, in which case we denote this minimum number by $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$. To calculate $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$, it follows from Proposition 4.3 that it is sufficient to consider a single extension of each agreement digraph for ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$. Referring back to Figure 1, observe that each of the three phylogenetic digraphs ${\mathcal{D}}_{1}$, ${\mathcal{D}}_{2}$, and ${\mathcal{D}}_{3}$ is in fact an agreement digraph for the two tree-child networks ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ that are shown in the same figure.
@@
SNPR± If $u$ is a tree vertex, then delete $e$, suppress $u$, subdivide an edge that is not a descendant of $v$ with a new vertex $u^{\prime}$, and add the new edge $(u^{\prime},v)$.
SNPR- If $u$ is a tree vertex and $v$ is a reticulation, then delete $e$, and suppress $u$ and $v$.
SNPR+ Subdivide $e$ with a new vertex $v^{\prime}$, subdivide an edge in the resulting network that is not a descendant of $v^{\prime}$ with a new vertex $u^{\prime}$, and add the new edge $(u^{\prime},v^{\prime})$.
By definition of a tree vertex, $u\neq\rho$ if we apply an SNPR±. If it is not important which of SNPR-, SNPR+, and SNPR± has been applied to ${\mathcal{N}}$ we simply refer to it as an $\mathrm{SNPR}$. By [6, Proposition 3.1], the operation is reversible, i.e. if ${\mathcal{N}}^{\prime}$ is a phylogenetic network on $X$ that can be obtained from ${\mathcal{N}}$ by a single SNPR, then ${\mathcal{N}}$ can also be obtained from ${\mathcal{N}}^{\prime}$ by a single SNPR. Lastly, we note that the well-known rSPR operation is an application of SNPR± to a phylogenetic tree.
@@
and refer to $w(\sigma)$ as the weight of $\sigma$. Intuitively, each deletion or addition of an edge contributes one to $w(\sigma)$. Thus, since an SNPR± deletes an edge and, subsequently adds a new edge, such an operation adds two to $w(\sigma)$. It was shown in [6, Proposition 3.2] that any two phylogenetic networks ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ on the same leaf set are connected by an $\mathrm{SNPR}$ sequence. If ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ are tree-child, then, by the same proposition, there also exists an $\mathrm{SNPR}$ sequence that connects ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ such that each network in the sequence is tree-child. We refer to such an SNPR sequence as a tree-child $\mathrm{SNPR}$ sequence. Moreover, we define the tree-child $\mathrm{SNPR}$ distance $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ between two tree-child networks ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ as the minimum weight of any tree-child $\mathrm{SNPR}$ sequence connecting ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$. If ${\mathcal{N}}$ and ${\mathcal{N}}^{\prime}$ are two phylogenetic $X$-trees, then $d_{\mathrm{rSPR}}({\mathcal{N}},{\mathcal{N}}^{\prime})$ denotes the minimum number of rSPR operations needed to transform ${\mathcal{N}}$ into ${\mathcal{N}}^{\prime}$.
Global assumption. Let $\sigma=({\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{t})$ be an SNPR sequence that connects the two phylogenetic networks ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$ on $X$. Throughout the remainder of the paper, we assume that there exists no $i\in\{1,2,\ldots,t\}$, such that ${\mathcal{N}}_{i}$ can be obtained from ${\mathcal{N}}_{i-1}$ by an SNPR± that deletes a reticulation edge in ${\mathcal{N}}_{i-1}$. Indeed, if such an $i$ exists, then there exists an SNPR sequence $\sigma^{\prime}=({\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{i-1},{\mathcal{N}}_{i}^{\prime},{\mathcal{N}}_{i},{\mathcal{N}}_{i+1},\ldots,{\mathcal{N}}_{t})$ such that ${\mathcal{N}}_{i}^{\prime}$ can be obtained from ${\mathcal{N}}_{i-1}$ by an SNPR- and ${\mathcal{N}}_{i}$ can be obtained from ${\mathcal{N}}_{i}^{\prime}$ by an SNPR+. Since we are interested in SNPR sequences of minimum weight and $w(\sigma^{\prime})=w(\sigma)$, no generality is lost.
6. Bounding the tree-child SNPR distance
In this section, we establish Theorem 1.1 and show that the bounds are tight. Let ${\mathcal{N}}$ and ${\mathcal{N}}’$ ’ be two tree-child networks, and let ${\mathcal{R}}$ be an extension of a tree-child digraph ${\mathcal{D}}$ of ${\mathcal{N}}$. We first show that, if ${\mathcal{N}}$, ${\mathcal{N}}’$ ’, and ${\mathcal{R}}$ satisfy certain properties, then there exists an extension of ${\mathcal{D}}$ in ${\mathcal{N}}^{\prime}$ such that the cut sizes of ${\mathcal{D}}$ in ${\mathcal{N}}$ and ${\mathcal{N}}’$ ’ differ by at most one.
Lemma 6.1.

### v2

@@ 4 Properties of extensions
    Suppose that both of $e$ and $e^{\prime}$ are contained in one of $\mathcal{M}$} and $\mathcal{M}$$^{\prime}$, and neither $e$ nor $e^{\prime}$ is contained in the other embedding. Without loss of generality, we may assume that both of $e$ and $e^{\prime}$ are contained in $\mathcal{M}$} and, therefore, in $\mathcal{R}$}. As $\mathcal{M}$$^{\prime}$ does not contain a vertex with out-degree zero that is not in $X\cup\{\rho\}$, it follows that $v\notin{\mathcal{M}}^{\prime}$. Thus, by Lemma 6(ii), exactly one of $e$ and $e^{\prime}$ is in $\mathcal{R}$$^{\prime}$.
@@
    Suppose that exactly one of $e$ and $e^{\prime}$ is contained in one of $\mathcal{M}$} and $\mathcal{M}$$^{\prime}$ and neither $e$ nor $e^{\prime}$ is contained in the other embedding. Without loss of generality, we may assume that $\mathcal{M}$} contains $e$ and does not contain $e^{\prime}$, and that $\mathcal{M}$$^{\prime}$ contains neither $e$ nor $e^{\prime}$. Evidently $e\in{\mathcal{R}}$ and, by the definition of an extension, $e^{\prime}\notin{\mathcal{R}}$. Moreover, as $v\notin{\mathcal{M}}^{\prime}$ it follows from Lemma 6(ii) that exactly one of $e$ and $e^{\prime}$ is in $\mathcal{R}$$^{\prime}$.
@@
    Suppose that exactly one of $e$ and $e^{\prime}$ is contained in each of $\mathcal{M}$} and $\mathcal{M}$$^{\prime}$. Again by Lemma 6(i), exactly one of $e$ and $e^{\prime}$ is contained in each of $\mathcal{R}$} and $\mathcal{R}$$^{\prime}$.
@@
    Suppose that both of $e$ and $e^{\prime}$ are contained in one of $\mathcal{M}$} and $\mathcal{M}$$^{\prime}$, and exactly one of $e$ and $e^{\prime}$ is contained in the other embedding. Without loss of generality, we may assume that both of $e$ and $e^{\prime}$ are contained in $\mathcal{M}$} and, therefore, in $\mathcal{R}$}. Then, by Lemma 6(i), exactly one of $e$ and $e^{\prime}$ is contained in $\mathcal{R}$$^{\prime}$.
@@
The last proposition motivates the following terminology. Let $\mathcal{D}$} be a phylogenetic digraph of a phylogenetic network $\mathcal{N}$} on $X$, and let $\mathcal{R}$} be an extension of $\mathcal{D}$} in $\mathcal{N}$}. We set $c_{\mathcal{D}}=|E_{\mathcal{N}}-E_{\mathcal{R}}|$ and refer to $c_{\mathcal{D}}$ as the cut size of $\mathcal{D}$} in $\mathcal{N}$}. By Proposition 7, $c_{\mathcal{D}}$ is well defined.
@@
Let $\mathcal{M}$} be the embedding of $\mathcal{D}$} in $\mathcal{N}$} that underlies $\mathcal{R}$}. If $\mathcal{R}$} is not a root extension of $\mathcal{D}$} in $\mathcal{N}$}, then there exists a reticulation edge $(u,v)$ in $\mathcal{N}$} such that $(u,v)\in{\mathcal{R}}$ and $(u,v)\notin{\mathcal{M}}$. Since $\mathcal{N}$} is tree-child, $u$ is a tree vertex. Let $v^{\prime}$ be the child of $u$ in $\mathcal{N}$} with $v\neq v^{\prime}$. Clearly, $v^{\prime}$ is a tree vertex or a leaf in $\mathcal{N}$}, and $(u,v^{\prime})\notin{\mathcal{R}}$. Let $R_{v}$ be the component in $\mathcal{R}$} that contains $v$, and let $R_{v^{\prime}}$ be the component in $\mathcal{R}$} that contains $v^{\prime}$. If $R_{v}$ and $R_{v^{\prime}}$ are distinct, obtain $R^{\prime}_{v}$ from $R_{v}$ by deleting each vertex $s$ for which there exists a directed path from $s$ to $u$, and obtain $R^{\prime}_{v^{\prime}}$ from $R_{v^{\prime}}$ by adding $(u,v^{\prime})$ and each edge of $R_{v}$ that is contained in a directed path ending at $u$ and, if $R_{v}$ and $R_{v^{\prime}}$ are not distinct, obtain $R^{\prime}_{v}$ from $R_{v}$ by deleting $(u,v)$ and adding $(u,v^{\prime})$. Then ${\mathcal{R}}^{\prime}=({\mathcal{R}}-\{R_{v},R_{v^{\prime}}\})\cup\{R_{v}^{\prime},R_{v^{\prime}}^{\prime}\}$ is an extension of $\mathcal{D}$} in $\mathcal{N}$} with cut size $c_{\mathcal{D}}$. It is straightforward to check that $\mathcal{M}$} is the embedding of $\mathcal{D}$} in $\mathcal{N}$} that underlies $\mathcal{R}$$^{\prime}$ and that $\mathcal{R}$$^{\prime}$ has one reticulation edge less than $\mathcal{R}$}. If $\mathcal{R}$$^{\prime}$ is not a root extension of $\mathcal{D}$} in $\mathcal{N}$}, then repeat the construction for a reticulation edge in $\mathcal{N}$} that is in $\mathcal{R}$$^{\prime}$ but not in $\mathcal{M}$} until no such edge remains. This completes the proof of the lemma. ∎
@@
Figure 5: A phylogenetic network $\mathcal{N}$}, a phylogenetic digraph $\mathcal{D}$} of $\mathcal{N}$}, and two root extensions $\mathcal{R}$} and $\mathcal{R}$$^{\prime}$ (indicated by red lines) of $\mathcal{D}$} in $\mathcal{N}$} with $|E_{\mathcal{N}}-E_{\mathcal{R}}|=6\neq 5=|E_{\mathcal{N}}-E_{{\mathcal{R}}^{\prime}}|$.
We end this section by noting that Proposition 7 does not hold for two root extensions of arbitrary phylogenetic networks (see Figure 5 for an example). Nevertheless, we have the following result for tree-child networks. Its proof is similar to that of Proposition 7 and is omitted.
Proposition 9.
@@
5 Measures of dissimilarity between two tree-child networks
@@
Let $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ be two tree-child networks on $X$. Let ${\mathcal{D}}=\{D_{\rho},D_{1},D_{2},\ldots,D_{k}\}$ be a collection of leaf-labelled acyclic digraphs. If $\mathcal{D}$} is a phylogenetic digraph of $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, we refer to $\mathcal{D}$} as an agreement digraph of $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$. Now let $\mathcal{D}$} be an agreement digraph for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$. If $\mathcal{D}$} is tree-child, we say that $\mathcal{D}$} is an agreement tree-child digraph for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$. In the remainder of this paper, we are particularly interested in agreement digraphs of $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ whose cut size is minimum. To this end, let $\mathcal{D}$} be an agreement tree-child digraph for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, and let $c_{\mathcal{D}}$ and $c^{\prime}_{\mathcal{D}}$ be the cut size of $\mathcal{D}$} in $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, respectively. Then $\mathcal{D}$} is called a maximum agreement tree-child digraph for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ if the sum $c_{\mathcal{D}}+c^{\prime}_{\mathcal{D}}$ is minimum over all agreement tree-child digraphs for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$, in which case $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ denotes this minimum number. To calculate $m_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$, it follows from Proposition 7 that it is sufficient to consider a single extension of each agreement digraph for $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$. Referring back to Figure 2, observe that each of the three phylogenetic digraphs ${\mathcal{D}}_{1}$, ${\mathcal{D}}_{2}$, and ${\mathcal{D}}_{3}$ is in fact an agreement digraph for the two tree-child networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ that are shown in the same figure.
@@
  1. SNPR± If $u$ is a tree vertex, then delete $e$, suppress $u$, subdivide an edge that is not a descendant of $v$ with a new vertex $u^{\prime}$, and add the new edge $(u^{\prime},v)$.
  2. SNPR- If $u$ is a tree vertex and $v$ is a reticulation, then delete $e$, and suppress $u$ and $v$.
  3. SNPR+ Subdivide $e$ with a new vertex $v^{\prime}$, subdivide an edge in the resulting network that is not a descendant of $v^{\prime}$ with a new vertex $u^{\prime}$, and add the new edge $(u^{\prime},v^{\prime})$.
By definition of a tree vertex, $u\neq\rho$ if we apply an SNPR±. If it is not important which of SNPR-, SNPR+, and SNPR± has been applied to $\mathcal{N}$} we simply refer to it as an $\mathrm{SNPR}$. As observed by Bordewich et al. [7, Proposition 3.1], the operation is reversible, i.e. if $\mathcal{N}$$^{\prime}$ is a phylogenetic network on $X$ that can be obtained from $\mathcal{N}$} by a single SNPR, then $\mathcal{N}$} can also be obtained from $\mathcal{N}$$^{\prime}$ by a single SNPR. Lastly, we note that the well-known rSPR operation is an application of SNPR± to a phylogenetic tree.
@@
and refer to $w(\sigma)$ as the weight of $\sigma$. Intuitively, each deletion or addition of an edge contributes one to $w(\sigma)$. Thus, since an SNPR± deletes an edge and, subsequently adds a new edge, such an operation adds two to $w(\sigma)$. It was shown by Bordewich et al. [7, Proposition 3.2] that any two phylogenetic networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ on the same leaf set are connected by an $\mathrm{SNPR}$ sequence. If $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ are tree-child, then, by the same proposition, there also exists an $\mathrm{SNPR}$ sequence that connects $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ such that each network in the sequence is tree-child. We refer to such an SNPR sequence as a tree-child $\mathrm{SNPR}$ sequence. Moreover, we define the tree-child $\mathrm{SNPR}$ distance $d_{\rm tc}({\mathcal{N}},{\mathcal{N}}^{\prime})$ between two tree-child networks $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ as the minimum weight of any tree-child $\mathrm{SNPR}$ sequence connecting $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$. If $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ are two phylogenetic $X$-trees, then $d_{\mathrm{rSPR}}({\mathcal{N}},{\mathcal{N}}^{\prime})$ denotes the minimum number of rSPR operations needed to transform $\mathcal{N}$} into $\mathcal{N}$$^{\prime}$.
Global assumption. Let $\sigma=({\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{t})$ be an SNPR sequence that connects the two phylogenetic networks ${\mathcal{N}}_{0}$ and ${\mathcal{N}}_{t}$ on $X$. Throughout the remainder of the paper, we assume that there exists no $i\in\{1,2,\ldots,t\}$, such that ${\mathcal{N}}_{i}$ can be obtained from ${\mathcal{N}}_{i-1}$ by an SNPR± that deletes a reticulation edge in ${\mathcal{N}}_{i-1}$. Indeed, if such an $i$ exists, then there exists an SNPR sequence
$\sigma^{\prime}=({\mathcal{N}}_{0},{\mathcal{N}}_{1},{\mathcal{N}}_{2},\ldots,{\mathcal{N}}_{i-1},{\mathcal{N}}_{i}^{\prime},{\mathcal{N}}_{i},{\mathcal{N}}_{i+1},\ldots,{\mathcal{N}}_{t})$
such that ${\mathcal{N}}_{i}^{\prime}$ can be obtained from ${\mathcal{N}}_{i-1}$ by an SNPR- and ${\mathcal{N}}_{i}$ can be obtained from ${\mathcal{N}}_{i}^{\prime}$ by an SNPR+. Since we are interested in SNPR sequences of minimum weight and $w(\sigma^{\prime})=w(\sigma)$, no generality is lost.
6 Bounding the tree-child SNPR distance
In this section, we establish Theorem 1 and show that the bounds are tight. Let $\mathcal{N}$} and $\mathcal{N}$}’ be two tree-child networks, and let $\mathcal{R}$} be an extension of a tree-child digraph $\mathcal{D}$} of $\mathcal{N}$}. We first show that, if $\mathcal{N}$}, $\mathcal{N}$}’, and $\mathcal{R}$} satisfy certain properties, then there exists an extension of $\mathcal{D}$} in $\mathcal{N}$$^{\prime}$ such that the cut sizes of $\mathcal{D}$} in $\mathcal{N}$} and $\mathcal{N}$}’ differ by at most one.
Lemma 10.